OpenStudy (anonymous):

lim (h->25) (sqrt(h)-5)/(h-25) is

OpenStudy (anonymous):

substitute h=25 into equation, (sqrt(25)-5)/(25-25) 0/0 undefined apply L'Hopital's Rule differentiate all ((1/sqrt(h))/1 substitute again (1/5)/1 1/5

OpenStudy (valpey):

I think G0lden is close but maybe: \[\frac{d(\sqrt(h)-5)}{dy} = \frac{1}{(2\sqrt(h))}\] \[(1/10)/1 = 1/10\]

OpenStudy (valpey):

And of course I meant \[\frac{d()}{dh}\]

OpenStudy (anonymous):

i got both of those answers for my answer choices so not sure which is right but thanks

OpenStudy (valpey):

G0lden gave me a medal so I think he agrees with me. Plus I'm right.